OpenStudy (anonymous):

Write a recursive formula for the sequence 7, 4, 1, –2, –5, . . . Then find the next term.

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

aN= aN-1 -3 for all N>1 given a1=7

OpenStudy (anonymous):

Thank You!

OpenStudy (kenljw):

Just looking, it looks like it a decreasing function by 3 each time starting with thererfore I sub (n+1) = I sub n -3 : I sub 0 = 7